OpenStudy (anonymous):

What is the equation of the line that passes through the point (4, 1) and has a slope of 5? y = 5x - 3 y = 5x - 1 y = 5x + 3 y = 5x - 19

OpenStudy (anonymous):

@Helpmeplease99

OpenStudy (rmrjr22):

d

OpenStudy (anonymous):

i aggree with rmrjr22 the last option.

OpenStudy (rmrjr22):

insert the points into each question to find out the x or y intercepts

OpenStudy (anonymous):

so if u sub in 4 for x it gives y = 1 as per the equation is that right?

OpenStudy (rmrjr22):

Bingo!

OpenStudy (rmrjr22):

making it true

OpenStudy (anonymous):

yep last option.

OpenStudy (jhannybean):

you want o find the equation of the line in slope intercept form. Slope Intercept Form: \(\large y=mx+b\) To find this equation we'll need a couple variables,or ingredients. 1. \(\large m\) = our slope: 5 2. \(\large b\) = y-intercept =unknown 3. our points \(\large (x,y)\) : (4,1) plug them all into the equation: \(\large 1 = 5(4)+b\) \(\large b=-19\) Now that we've found all out variables,plug it back into the equation \(\large y=5x-19\)

OpenStudy (rmrjr22):

just got warned...